OpenStudy (anonymous):

solve 5x^2+12=-6x using quadratic formula

OpenStudy (anonymous):

Rewrite it into standard form: 5x^2+6x+12=0, you can use the quadratic formula from here

OpenStudy (anonymous):

\[x=-b \pm \sqrt{b^2-4ac}\] over 2a

OpenStudy (anonymous):

\[\frac{-6 \pm \sqrt{6^2-4(5)(12)}}{2(5)}\] Both solutions will be complex numbers

OpenStudy (anonymous):

\[\frac{-6 \pm \sqrt{-204}}{10}\] \[\frac{-6 \pm i \sqrt{204}}{10} = \frac{-6 \pm 2i \sqrt{51}}{10}\] =\[\frac{-3 \pm i \sqrt{51}}{5}\]

OpenStudy (anonymous):

what is the i for?

OpenStudy (anonymous):

Are you unfamiliar with imaginary numbers?

OpenStudy (anonymous):

i^2=-1

OpenStudy (anonymous):

no never used them

OpenStudy (anonymous):

Well, there are no REAL NUMBER solutions to the quadratic, but there are complex solutions

OpenStudy (anonymous):

and how do i do that

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Complex_number

OpenStudy (anonymous):

If you aren't expected to know imaginary/complex numbers, then I suppose you can write no solution...

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Imaginary_numbers

OpenStudy (anonymous):

thanks